OpenStudy (mkmkasim):

Help Pls! ∫(sinx)^4 dx

OpenStudy (turingtest):

you can either use the reduction formulas, or repeatedly apply the identity\[\sin^2x=\frac12[1-\cos(2x)]\]

OpenStudy (turingtest):

followed by\[\cos^2u=\frac12[1+\cos(2u)]\]

OpenStudy (mkmkasim):

well I broke it down like this, ∫(sinx)^4=∫((sinx)^2)^2=∫((1+cos2x)/2)^2

OpenStudy (mkmkasim):

then I got stuck...

OpenStudy (turingtest):

expand that and what do you get?

OpenStudy (turingtest):

wait wait, first of all that should be 1-cos(2x)

OpenStudy (mkmkasim):

i belive it is 1/4∫1-2cos2x+(cosx)^2

OpenStudy (turingtest):

the last term is wrong, why did you change the argument from cos(2x) to cosx?

OpenStudy (mkmkasim):

right right it should be (cos2x)^2 correct?

OpenStudy (turingtest):

right

OpenStudy (turingtest):

so that term needs to be dealt with to be able to integrate it, so what formula can we apply here?

OpenStudy (mkmkasim):

im drawing a blank here...

OpenStudy (mkmkasim):

this is where i got stuck

OpenStudy (turingtest):

we only have two to work with\[\sin^2u=\frac12[1-\cos(2u)]\]\[\cos^2u=\frac12[1+\cos(2u)]\]

OpenStudy (turingtest):

"but wait" you say, "I have cos(2x), not cos(x), how can I apply the second formula?" answer: let u=2x

OpenStudy (mkmkasim):

ahhh....

OpenStudy (mkmkasim):

so 1/8∫1-2cosu+(cosu)^2 ?

OpenStudy (turingtest):

yeah, now how would you change (cosu)^2 ?

OpenStudy (mkmkasim):

since you have cosu already would you change it to 1-(sinu)^2?

OpenStudy (turingtest):

no, because how the heck do you integrate sin^2 ? we want to get everything to the power of 1, so apply one of the double angle formulas I have written above

OpenStudy (mkmkasim):

so (1+cos2u)/2

OpenStudy (turingtest):

correct, now you can re-substitute 2x for u

OpenStudy (mkmkasim):

i think i may have messed this up but is it 1+cos4x/2 ?

OpenStudy (turingtest):

you only messed up in the lack of parentheses (1+cos4x)/2 ?

OpenStudy (mkmkasim):

right...

OpenStudy (turingtest):

so now the whole thing is....?

OpenStudy (mkmkasim):

1/16∫1-2cos2x+1+cos4x

OpenStudy (turingtest):

you didn't simplify right 1/4∫1-2cos2x+(cosx)^2dx= 1/4∫1-2cos2x+(1+cos4x)/2dx=1/8∫2-4cos2x+1+cos4xdx=...

OpenStudy (mkmkasim):

ok so the 1/2 is the same as multiplying the other terms by 2 then...

OpenStudy (turingtest):

yes, if you factor 1/2 out of the integrand, then you factor 1/2 out of each term, which is the same as dividing each term by 1/2, and x/(1/2)=2x

OpenStudy (mkmkasim):

ok, so now you can just push the integral inside and solve for each one then correct?

OpenStudy (mkmkasim):

keeping the 1/8 with the integral for each one...

OpenStudy (turingtest):

yep, now integrate term-by-term I'd keep the 1/8 outside until the very end

OpenStudy (mkmkasim):

ok i think i got it thanks...

OpenStudy (turingtest):

welcome !

OpenStudy (mkmkasim):

just curious are you an undergrad or graduate student?

OpenStudy (turingtest):

I am neither. Completely self-taught, basically. Never been to college, so feel free to doubt me on anything :D

OpenStudy (mkmkasim):

oh haha. well good job teaching yourself!

OpenStudy (turingtest):

Thanks, I just happen to like this stuff. I hope to go to school eventually...